OpenStudy (anonymous):

22. At what angle with the horizontal should a hunter aim his gun so he will hit a target is 60 meters horizontally away from him, and 20.0 mm above him, if the muzzle velocity of the bullet is 400 m/s? got 81.4 deg. right?

OpenStudy (anonymous):

mm should be in meters i suppose

OpenStudy (anonymous):

 let angle with horizontal be X the verticle component of velocity is u*sinX and horizontal componet of velocity is u*cosX. let after time T it will hit the target then distance between hunter and target=u*cosX*T find the value of T from this equation i.e T=60/(u*cosX). and in the same time it also travelles the vertical distance(distance between target and ground).i.e we apply the basic formula d=usinX*T-1/2gT^2 put the value of d=20m,u=400 and T=60/u*cosX. and you get an equation in one variable that variable is X hence after solving it ,get your answer,if u have any problem ,talk to me.